log x + log y (x-21)=2
Answer by lwsshak3(6505)

solve for x
log x + log y (x-21)=2
log[x(x-21)]=2
convert to exponential form: base(10) raised to log of number(2)=number(x(x-21))
10^2=x^2-21x
x^2-21x-100=0
(x-25)(x+4)=0
x=25
or
x=-4